OpenStudy (anonymous):

(x+3)^5 using pascal's triangle

OpenStudy (anonymous):

To check your answer take a look at http://www.wolframalpha.com/input/?i=expand+%28x%2B3%29^5

OpenStudy (anonymous):

1 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1

OpenStudy (anonymous):

\[ (x+3)^5 = 1 x^5 + 5 x^4(3) + 10 x^3 (3^2) + 10 x^2 (3^3)+ 5 x (3)^4 + 3^5 \]

OpenStudy (anonymous):

\[ (x+3)^5=x^5+15 x^4+90 x^3+270 x^2+405 x+243 \]